OpenStudy (anonymous):

If 1/2 an apple was shared equally among the 3 people, what fraction of the whole apple did each person receive? A) 1 6 B) 3 6 C) 5 12 D) 5 16

OpenStudy (anonymous):

how would I find the answer

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

You divide 1/2 / 3

OpenStudy (igreen):

Or multiply by the reciprocal: \(\dfrac{1}{2} \times \dfrac{1}{3}\) = ?

OpenStudy (igreen):

Multiply the tops together and the bottoms.

OpenStudy (anonymous):

\(\dfrac{1}{2}{ \div\ 3}\)

OpenStudy (igreen):

Tell me what you get.

OpenStudy (anonymous):

a 1/6

OpenStudy (anonymous):

3/6

OpenStudy (anonymous):

LOL no.

OpenStudy (anonymous):

3/6 is totally wrong.

OpenStudy (igreen):

Yes, A is correct.